# Enhancing Data-Limited Graph Neural Networks by Actively Distilling Knowledge from Large Language Models

### v1

@@ 5.3. Evaluation of Hyper-parameters
We evaluate the impact of different hyper-parameters to answer RQ2. We evaluate our proposed model with different hyper-parameter settings: training size $\mathbf{N}\in\{C\times 1,C\times 3,C\times 5,C\times 7\}$; the budget size $B\in[1,2,3,4,5,7,10,15,20,25]$ for graph-LLM based active learning with $N=C\times 3$; the balance parameters $\alpha\in(0,0.5]$ with $\beta=0.1$ and $\beta\in[0.01,0.03,0.05,0.1,0.2,0.3,0.4,0.5]$ with $\alpha=0.3$, respectively; and the KD temperature $\tau\in[1,5]$. The evaluation results for training size are shown in Tables 3, 4, and 5. The hyper-parameter evaluations are illustrated in Figure 3.
@@
  • From Table 3, Table 4, and Table 5, we can easily observe that when we increase the training size $\mathbf{N}$, the performance in terms of classification accuracy consistently improves. Especially when $\mathbf{N}$ is increased from $1$ to $3$, the performance significantly improves.
@@
  • In Figure 3(a), we observe that as the value of $\alpha$ increases, the performance initially improves, and then reaches a peak around $\alpha=0.3$. However, beyond this optimal point, the performance experiences a sharp decline with further increases in $\alpha$.
@@
Figure 3. Hyper-parameters evaluation: (a) Alpha (b) Beta (c) Temperature (d) Budget Size
5.4. Ablation Study
@@
Table 6. Ablation study: Soft Labels (SL), Enhanced Rationales (ER), and Graph-LLM-based active learning (AL)
SL ER AL Citeseer Cora PubMed
51.51±(3.88) 65.90±(7.06) 62.74±(1.41)
53.00±(1.43) 68.87±(4.82) 62.23±(1.20)
54.39±(3.84) 70.32±(3.81) 63.82±(0.99)
54.65±(2.58) 67.99±(3.12) 63.19±(2.41)
55.57±(2.18) 71.12±(2.63) 63.62±(1.21)
55.41±(1.13) 72.91±(1.74) 63.96±(0.89)
54.49±(2.93) 71.78±(2.80) 63.68±(0.82)
57.63±(3.58) 74.30±(2.49) 65.01±(0.73)
As illustrated in Table 6, all these components contribute to the performance of our model. Among these components, the enhanced rationales have a relatively small impact on the performance. When we add soft-labels and active learning independently, the performance improves by a considerable margin. When we add these two components together, the performance has a significant improvement. The experimental results also showcase that LLMs can effectively enhance the performance of GNNs. Whether we incorporate the logits and enhanced rationales independently or combine them, there is a significant performance improvement. Furthermore, when all these components are integrated, our model achieves state-of-the-art performance.
For the rationale alignments, we further evaluate the two different alignment strategies: 1) max pooling and 2) our MLP-based alignment approach. For active learning, we evaluate different selection strategies: 1) randomly select nodes with pseudo-labels 2). select all valuable nodes at once with Graph-LLM-based AL; 2) select nodes in an iteration method with our Graph-LLM-based AL, which means we will select $b$ nodes per class until the total number of selected nodes reach $B\times C$. As shown in Table 7, compared with the MLP-based alignment strategy, the improvement by using max pooling is lower, which is reasonable because max pooling will lose some information during the pooling process. For the active learning strategy, the performance of the model with our Graph-LLM-based AL is better than selecting nodes randomly. Moreover, despite selecting all valuable nodes at once has shown significant performance improvement, the performance reaches a new high when we apply the iteration selection strategy. This enhancement is attributed to the iterative active learning process, where GNNs benefit from the LLM’s zero-shot inference and reasoning ability to refine their predictions iteratively.
Table 7. Ablation study: alignment and active learning
@@
Alignment: max pooling 52.31 ±(2.38) 71.64±(3.44) 62.51±(1.44)
AL: Random Selection 55.79±(2.07) 72.77±(1.96) 63.63±(0.61)
AL: w/o iteration 55.96±(3.16) 73.01±(2.13) 64.76±(0.57)
Our Model 57.63±(3.58) 74.30±(2.49) 65.01±(0.73)
6. Conclusion
In this paper, we extend the task of node classification to a more challenging and realistic case where only few labeled data are available. To tackle this challenge, we propose a novel few-shot node classification model that leverages the zero-shot and reasoning ability of Large Language Models. We treat LLMs as a teacher to teach GNNs from two different perspectives including the logits from the distribution side and enhanced rationales from the feature side. Moreover, we proposed a Graph-LLM-based active learning method to further improve the generalizability of GNNs with few available data by actively selecting and distilling knowledge from LLMs. To assess the effectiveness of our model, extensive experiments have been conducted on three citation networks. The evaluation results demonstrate that our model achieves state-of-the-art performance and LLMs can effectively provide insights to GNNs from different perspectives, reaffirming its effectiveness in node classification, its superiority over baseline methods, and its practical significance in addressing the challenges of few-shot node classification.

### v2

@@ V-C Hyper-parameter Sensitivity Analysis
We evaluate the impact of different hyper-parameters to answer RQ2. We evaluate our proposed model with different hyper-parameter settings: training size $\mathbf{N}\in\{C\times 1,C\times 3,C\times 5,C\times 7\}$; the budget size $B\in[1,2,3,4,5,7,10,15,20,25]$ for graph-LLM based active learning with $N=C\times 3$; the balance parameters $\alpha\in(0,0.5]$ with $\beta=0.1$ and $\beta\in[0.01,0.03,0.05,0.1,0.2,0.3,0.4,0.5]$ with $\alpha=0.3$, respectively; and the KD temperature $\tau\in[1,5]$. The evaluation results for training size are shown in Tables IV, III, and V. The hyper-parameter evaluations are illustrated in Figure 3.
@@
  • From Table IV, Table III, and Table V, we can easily observe that when we increase the training size $\mathbf{N}$, the performance in terms of classification accuracy consistently improves. Especially when $\mathbf{N}$ is increased from $1$ to $3$, the performance significantly improves.
@@
  • In Figure 3(a), we observe that as the value of $\alpha$ increases, the performance initially improves, and then reaches a peak around $\alpha=0.3$. However, the performance decreases drastically with further increases in $\alpha$ beyond this point.
@@
Figure 3: Hyper-parameters evaluation: (a) Alpha (b) Beta (c) Temperature (d) Budget Size
TABLE VI: Ablation study: Soft Labels (SL), Enhanced Rationales (ER), and Graph-LLM-based active learning (AL)
SL ER AL Citeseer Cora PubMed
52.24±(1.04) 66.27±(5.04) 62.64±(0.85)
53.21±(2.25) 68.93±(3.88) 63.23±(1.15)
54.39±(3.84) 70.32±(3.81) 63.82±(0.99)
54.65±(2.58) 68.87±(2.12) 63.19±(2.41)
55.57±(2.18) 71.12±(2.63) 64.33±(1.27)
55.41±(1.13) 72.91±(1.74) 64.59±(1.02)
55.37±(1.83) 71.78±(2.80) 63.68±(0.82)
57.43±(3.33) 74.32±(1.79) 65.40±(0.84)
V-D Ablation Study
@@
TABLE VII: Ablation study: alignment and active learning
@@
Alignment: max pooling 52.28±(2.37) 70.88±(2.24) 62.43±(1.21)
AL: Random Selection 54.97±(2.32) 72.77±(1.86) 63.58±(0.64)
AL: w/o iteration 55.84±(2.83) 73.12±(1.97) 64.43±(0.57)
Our Model 57.43±(3.33) 74.32±(1.79) 65.40±(0.84)
As illustrated in Table VI, all these components contribute to the performance of our model. Among these components, the enhanced rationales have a relatively small impact on the performance. When we add soft-labels and active learning independently, the performance improves by a considerable margin. When we add these two components together, the performance has a significant improvement. The experimental results also showcase that LLMs can effectively enhance the performance of GNNs. Whether we incorporate the logits and enhanced rationales independently or combine them, there is a significant performance improvement. Furthermore, when all these components are integrated, our model achieves state-of-the-art performance.
For the rationale alignments, we further evaluate the two different alignment strategies: 1) max pooling and 2) our MLP-based alignment approach. For active learning, we evaluate different selection strategies: 1) randomly select nodes with pseudo-labels 2). select all valuable nodes at once with Graph-LLM-based AL; 2) select nodes in an iteration method with our Graph-LLM-based AL, which means we will select $b$ nodes per class until the total number of selected nodes reach $B\times C$. As shown in Table VII, compared with the MLP-based alignment strategy, the improvement by using max pooling is limited, which is reasonable because max pooling will lose information during the pooling process. For active learning, the performance of the model with our Graph-LLM-based AL is better than random selection. Moreover, despite selecting all valuable nodes at once has shown significant performance improvement, the performance reaches a new high when we apply the iteration selection strategy. This enhancement is attributed to the iterative active learning process, where GNNs benefit from the LLM’s zero-shot inference and reasoning ability to refine their predictions iteratively.
VI Conclusion
In this paper, we extend the task of node classification to a more challenging and realistic case where only few labeled data are available. To tackle this challenge, we propose a novel few-shot node classification model that leverages the zero-shot and reasoning ability of Large Language Models. We treat LLMs as a teacher to teach GNNs from two different perspectives including the logits from the distribution side and enhanced rationales from the feature side. Moreover, we proposed a Graph-LLM-based active learning method to further improve the generalizability of GNNs with few available data by actively selecting and distilling knowledge from LLMs. To assess the effectiveness of our model, extensive experiments have been conducted on three citation networks. The evaluation results demonstrate that our model achieves state-of-the-art performance and LLMs can effectively provide insights to GNNs from different perspectives, reaffirming its effectiveness in node classification, its superiority over baseline models, and its practical significance in addressing the challenges of few-shot node classification.